OpenStudy (anonymous):

Mona has an account with a balance of $738. She originally opened the account with a $500 deposit and a simple interest rate of 5.6%. If there were no deposits or withdrawals, how long ago was the account opened?

OpenStudy (anonymous):

8.5 years

OpenStudy (anonymous):

can u explain how u got it

OpenStudy (anonymous):

p = 500 si = 738 - 500 = 238 r = 5.6 si = ptr/100 238 = 500*t*5.6/100 t=8.5

OpenStudy (anonymous):

do you understand, still you have doubt

OpenStudy (anonymous):

i understand thanks